# MemoFurther Readings

### v1

@@ Introduction
Coordinated actors may employ a variety of tactics to achieve their goals. A common tactic is to artificially amplify content on specific topics to create the appearance of widespread grassroots support and manipulate platforms’ feed algorithms (Mannocci et al., 2024; Vishnuprasad et al., 2024; Pacheco et al., 2020). To uncover content amplification, we construct a Text Similarity Network (TSN), where nodes represent users linked by the similarity of the content they share. This TSN is then employed to identify a subset of users exhibiting suspiciously coordinated behavior.
@@
The initial step in constructing the TSN involves preprocessing the raw data to ensure the results are meaningful. In line with standard practices (Luceri et al., 2024a), we exclude retweets and remove punctuation, stopwords, emojis, URLs, as well as any content with fewer than four words. To capture the semantic nuances of the content, we embed all text data using the SentenceTransformer model stsb-xlm-r-multilingual [^3] https://huggingface.co/sentence-transformers/stsb-xlm-r-multilingual, and then calculate the average cosine similarity between pairs of users. To avoid spurious correlations, we consider text similarity valid only if it occurs within a one-day sliding window. For each window, we calculate the text similarity between the shared content of user pairs, setting the edge weight in the TSN to the average similarity observed across all windows.
@@
To identify coordinated accounts, we apply established thresholds from the literature (Pacheco et al., 2021; Luceri et al., 2024a). Specifically, we filter out edges with similarity below 0.95 and classify users as coordinated if their nodes rank in the top 0.5% by eigenvector centrality.
Characterizing Content Pushed by CoIA
@@
In this section, we present an overview of the methods used to characterize textual content amplified by coordinated actors across various social media networks. This characterization encompasses topic analysis, AI-generated content detection, and an assessment of content credibility.
Topic Analysis
@@
Coordinated accounts often amplify specific narratives or themes to steer public attention toward polarized, inflammatory, or misleading discussions. To uncover the agendas these accounts seek to promote, we use BERTopic (Grootendorst, 2022), a state-of-the-art tool for topic extraction. For a detailed description of BERTopic’s methodology, we refer readers to (Grootendorst, 2022). This approach also helps identify patterns in shared content, improving our understanding of coordinated activities and their intended impact on public discourse.
Once coordinated accounts are identified using the methodology outlined in Sec. Detection of Content Amplification, we apply BERTopic to the entire content corpus. This allows us to map the topics promoted by coordinated accounts and compare them with those shared by organic users (i.e., users not classified as coordinated). We choose BERTopic over alternatives such as LDA or GPT-based approaches because it offers an effective balance between accuracy and scalability.
AI-Generated Content Detection
@@
To identify AI-generated text, we use the approach introduced by (Dmonte et al., 2024), which is designed to identify AI-generated content on $\mathbb{X}$. We initially trained a RoBERTa [^4] https://huggingface.co/FacebookAI/roberta-base model on their $\mathbb{X}$ dataset and then evaluated its performance on datasets from other platforms by developing a new validation set containing approximately 2,000 samples. This validation set was constructed using Llama 3.1 3B Instruct [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, GPT-4o [^6] https://openai.com/index/hello-gpt-4o/, and open-source social media datasets from Telegram, $\mathbb{X}$, and Facebook, sampled from posts between 2010 and 2015. We assume these older datasets do not contain AI-generated content and sampled around 1,000 sentences from them. To create a balanced set of human and AI-generated content, we prompted the language models to generate 1,000 samples that mirrored the original texts in topic distribution and length. This approach ensured that the AI-generated texts were comparable to human-generated samples in both thematic and structural characteristics. Results showed precision values ranging from 0.87 to 0.97 in detecting AI-generated content within the validation set, aligning with our goal of prioritizing precision over recall to minimize false positives (misclassified legitimate users). Full results are available in the Appendix.
Credibility Assessment
@@
We assess the credibility of web domains using Media Bias/Fact Check (MBFC). MBFC is an independent watchdog that rates news outlets on a 6-point factuality scale, ranging from Very Low to Very High. For each post in our datasets, we systematically extract, expand, and parse all embedded URLs, checking whether the URL belongs to a low- or high-credibility domain from our lists.
@@
Given recent news [^7] https://www.bbc.com/news/articles/c8rx28v1vpro [^8] https://www.state.gov/u-s-department-of-state-takes-actions-to-counter-russian-influence-and-interference-in-u-s-elections/ and prior instances of interference by Russian agencies in U.S. elections (Badawy et al., 2019), we also examined the potential amplification of Russian media outlets. Following a similar approach to previous work (Pierri et al., 2023), we utilize the VoynaSlov dataset (Park et al., 2022) to obtain a list of 23 state-affiliated Russian websites. We then check whether the extracted URLs link to one of these web domains.
@@
Finally, we employ a similar approach to identify content linked to conspiracies, such as QAnon, given the relevance of fringe theories during the 2020 US Election, and its aftermath (Sharma et al., 2022; Vishnuprasad et al., 2024). Building on the methodology of (Wang et al., 2023), we detect posts sharing QAnon content by utilizing a list of keywords commonly associated with the conspiracy, as outlined by (Sharma et al., 2022).
Results
@@
This section presents the results obtained from the detection of online coordination. We first present the results for the detection of web domain promotion and content amplification. Following the identification of these campaigns, we examine the content pushed by coordinated networks and the engagement received by suspicious actors focusing on several interaction and engagement metrics.
Detecting CoIA for Web Domain Promotion (RQ1)
@@
We detect CoIAs that push web domains within and across platforms as described in Section Detection of Web Domain Promotion. We begin by presenting results for each coordinated network within individual platforms. Following this, we analyze the inter-platform CoIA.
@@
Figure 1. Cross-platform coordination network showing user coordination across four social media networks.
Intra-Platform Coordination.
We first analyze each intra-platform CoIA based on the co-URL similarity networks extracted from each platform separately. We describe intra-platform CoIAs as follows.
Telegram: We identified 33 highly coordinated channels co-sharing URLs to web domains with a partisan slant and low factuality. As shown in the Table 2, the most frequently shared domains are predominantly right-leaning and of low credibility, as assessed by MBFC [^9] https://mediabiasfactcheck.com/. Notably, the table highlights the presence of a Russian state-controlled media outlet (RT.com) and its video-on-demand subsidiary (Ruptly.tv), which could indicate a foreign effort to interfere in the election. Both RT.com and Ruptly.tv have previously been accused of orchestrating campaigns to influence U.S. elections via social media [^10] https://www.nbcnews.com/politics/2020-election/facebook-blocks-russia-backed-accounts-other-sites-keep-churning-out-n1242683.
Additionally, this CoIA promotes a far-right website (thegatewaypundit.com) and an extremist-friendly video platform (odysee.com), suggesting ties to fringe ideas. A manual review of these Telegram channels, including their profile descriptions (see Table 8) and shared messages (see Table 12), reveals a strong prevalence of content and accounts promoting conspiracy theories, particularly those related to COVID vaccines, 5G, and alternative news media.
@@
Table 2. Top-5 domains shared in Telegram by coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on data from MBFC.
$\mathbb{X}$/Twitter: We identified a network of 19 coordinated accounts predominantly promoting content from right-leaning domains, as listed in Table 3. A manual review of this Twitter CoIA reveals that many of these coordinated accounts share similar profile descriptions, reflecting narratives tied to religious and conservative principles (see Table 9). Additionally, an analysis of the messages (see Table 13) shows a strong presence of politically partisan content, predominantly aligned with right-leaning ideologies.
@@
newsbreakapp.com 5 NA NA
go.shr.lc 2 NA NA
@@
Table 3. Top-5 domains shared in Twitter by coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on data from MBFC.
@@
Facebook: On Facebook, no well-known media outlets dominate the shared content. However, the most shared domain is a partisan news outlet, as suggested by the semantics of its title: gorightnews.com. Additionally, there is a direct connection to a public activist known for political campaigns [^11] https://en.wikipedia.org/wiki/Peter_Boykin and their website: peterboykin.com. These campaigns and related Facebook accounts primarily focus on the “Gays for Trump" movement, an American LGBTQ organization that advocates for former U.S. President Donald Trump and his administration  [^12] https://en.wikipedia.org/wiki/Gays_for_Trump. The top-engagement messages (see Table 14) and bios (see Table 10) indicate strong support for right-leaning ideologies.
Cross-Platform Coordination.
@@
The analysis of the cross-platform network reveals distinct patterns. The unified cross-platform network, displayed in Fig. 1, highlights a giant component connecting coordinated accounts from both Telegram and Twitter, while the Facebook coordination network remains largely disconnected. A few noteworthy observations:
@@
First, we observe that most cross-platform connections occur between Twitter accounts and Telegram channels, likely due to their higher representation in the dataset. Examining the bios of the top users by degree, as shown in the Appendix (see Table 15), we note a prominent presence of non-mainstream news outlets and conspiracy theories, such as the flat earth one.
@@
Second, when focusing on the bridge nodes between the Telegram-coordinated cluster and the cross-platform giant component, we find highly influential channels with up to 24,000 subscribers. These channels often feature bios referencing free speech and religion, such as:
  • "The FIGHT for FREEDOM - :Aron TRUTH Social"
  • "Q Reee-searchers Watchers"
@@
  • "’Arise and shine, for your light has come, and the glory of the LORD rises upon you.’ Isa. 60:1 WE RISE beyond the challenges of yesterday to become better stewards of tomorrow!"
@@
Similarly, the bridge node between the Twitter-coordinated cluster and the cross-platform giant component is a highly influential account with 32,000 followers. The bio reads: "NO DM’S. Beautiful disaster. Self-proclaimed arbiter of great ideas. Here to annoy the dumb asses. NO LISTS!! #Imvotingforafelon #animallover", indicating a strong partisan flavor.
@@
Finally, the largest component of the cross-platform coordination network is dominated by a mix of Telegram and Twitter accounts promoting domains such as magapac.com, QAnon-related narratives (e.g., "WWG1WGA" [^13] https://en.wiktionary.org/wiki/WWG1WGA), and accounts with partisan bios like:
@@
  • "We The Ultra Patriots is made up of patriotic Americans dedicated to exposing crimes against humanity, false flags, lies, and corruption. We The Ultra Patriots are…"
@@
A closer look at the top shared domains within the coordinated network (see Table 4) reveals that most are partisan, including truthsocial.com, an alternative non-mainstream social media platform [^14] https://en.wikipedia.org/wiki/Truth_Social. Additionally, seven of the shared domains are Russian state-affiliated websites [^15] https://github.com/chan0park/VoynaSlov/tree/master, including tv5, rbc, gazeta, ruptly, and mil, further suggesting potential Russian information operations to interfere with the election discourse. Examining the posts and messages with the highest engagement (see Table 15 in the Appendix), we observe the prevalence of ultra-MAGA narratives, conservative religious themes, and environmental news.
@@
Table 4. Top domains shared by cross-platform coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on data from MBFC.
@@
Figure 2. Content analysis of coordinated channels on Telegram, Twitter, and Facebook. We report the top-10 most recurring topics and compare their prevalence against the organic discourse.
Detecting CoIA for Content Amplification (RQ2)
In this section, we present our findings on the specific topics amplified by coordinated users within content amplification campaigns. Following the method in Sec. Detection of Content Amplification, we extract the Text Similarity Network and identify a seed set of coordinated users. We then apply BERTopic, as described in Sec. Topic Analysis, to identify the themes these users aim to promote and compare them to the themes that emerge organically on each platform. For each platform, we report the top-10 most prevalent topics within the coordinated cohort and compare their prevalence to that within the organic population.
This approach highlights themes promoted by coordinated users, with large differences indicating topics characteristic of coordinated activity. In the remainder of this section, we detail our findings on coordination-driven content amplification across the three platforms analyzed.
Telegram: Based on the significantly high similarity of their shared content, we identified 57 coordinated Telegram channels. Among these channels, public health is a prominent topic of discussion, as can be seen in Figure 2(a), with skepticism around the COVID-19 vaccine emerging as the most frequently discussed theme. Additionally, many messages promote alternative medicine, and there is notable discussion surrounding a conspiracy theory related to bird flu vaccination. Telegram’s coordinated channels also focus on strictly political discussions, particularly on the U.S. presidential debate, the Russia-Ukraine conflict, and Donald Trump’s legal challenges. Interestingly, we observe discourse around immigration issues in Ireland, which appears to be echoed by accounts associated with the MAGA movement. Notably, three out of the ten topics involve conspiracy theories related to flat-earth beliefs, bird flu, and antisemitism. The most representative content for each topic, as identified by BERTopic, is provided in the Appendix.
We also assess the presence of QAnon-related keywords in content shared by both coordinated and organic channels. Approximately 1.66% of content from coordinated users contains at least one QAnon-related keyword, compared to only 0.12% of content from the organic population. This substantial difference is evident in Table 5, where we list the top five keywords present in the coordinated group, also compared to keywords shared by organic channels. Although coordinated channels make up only 0.36% of the total Telegram dataset, they show a marked prominence, both in absolute and relative terms, in sharing QAnon-related keywords compared to the organic channels. These findings indicate a propensity for Telegram CoIAs to disseminate conspiracies.

### v2

@@ 4.1.2. Detection of Content Amplification
Coordinated actors may employ a variety of tactics to achieve their goals. A common tactic is to artificially amplify content on specific topics to create the appearance of widespread grassroots support and manipulate platforms’ feed algorithms (Mannocci et al., 2024; Vishnuprasad et al., 2024; Pacheco et al., 2020). To uncover content amplification, we construct a Text Similarity Network (TSN), where nodes represent users linked by the similarity of the content they share. This TSN is then employed to identify a subset of users exhibiting suspiciously high similarity, potentially indicating coordinated behavior (Pacheco et al., 2021).
@@
The initial step in constructing the TSN involves preprocessing the raw data to ensure the results are meaningful. In line with standard practices (Luceri et al., 2024b), we exclude retweets and remove punctuation, stopwords, emojis, URLs, as well as any content with fewer than four words. To capture the semantic nuances of the content, we embed all text data using the SentenceTransformer model stsb-xlm-r-multilingual555huggingface.co/sentence-transformers/stsb-xlm-r-multilingual, and then calculate the average cosine similarity between pairs of users. To avoid spurious correlations, we consider text similarity valid only if it occurs within a one-day sliding window. For each window, we calculate the text similarity between the shared content of user pairs, setting the edge weight in the TSN to the average similarity observed across all windows.
@@
To identify coordinated accounts, we apply established thresholds from the literature (Pacheco et al., 2021; Luceri et al., 2024b). Specifically, we filter out edges with similarity below 0.95 and classify users as coordinated if their nodes rank in the top 0.5% by eigenvector centrality. Due to substantial variability in textual content sharing, driven by platform-specific constraints, such as character limits on some platforms, and differences in the nature of messaging (e.g., Telegram chats versus Facebook and $\mathbb{X}$ posts), our analysis of content amplification is focused only on intra-platform CoIA.
4.2. Characterizing Content Pushed by CoIA
@@
In this section, we outline the methods used to characterize textual content amplified by coordinated actors. This characterization encompasses topic analysis, AI-generated content detection, and an assessment of content credibility.
4.2.1. Topic Analysis
@@
Coordinated accounts often amplify specific narratives or themes to steer public attention toward polarized, inflammatory, or misleading discussions. To uncover the agendas these accounts seek to promote, we use BERTopic (Grootendorst, 2022), a state-of-the-art tool for topic extraction. This approach also helps identify patterns in shared content, improving our understanding of coordinated activities and their intended impact on public discourse. Once coordinated accounts are identified using the methodology outlined in §4.1.2, we apply BERTopic to the entire content corpus. This allows us to map the topics promoted by coordinated accounts and compare them with those shared by organic users (i.e., users not classified as coordinated). We choose BERTopic over alternatives such as LDA or GPT-based approaches because it offers an effective balance between accuracy and scalability.
4.2.2. AI-Generated Content Detection
@@
AIGC has been linked to nefarious activities (Ferrara, 2024), such as orchestrating malicious campaigns. To identify AI-generated text, we use the approach introduced by (Dmonte et al., 2024), which is designed to identify AI-generated content on $\mathbb{X}$. We initially trained a RoBERTa666huggingface.co/FacebookAI/roberta-base model on their $\mathbb{X}$ dataset and then evaluated its performance on datasets from other platforms by developing a new validation set containing approximately 2,000 samples. Results showed precision values ranging from 0.87 to 0.97 in detecting AI-generated content within the validation set, aligning with our goal of prioritizing precision over recall to minimize false positives (i.e., misclassified legitimate users). A detailed description of how we built the validation set and the full results are available in the Appendix.
4.2.3. Credibility Assessment
@@
We assess the credibility of web domains using Media Bias/Fact Check (MBFC)—an independent watchdog that rates news outlets on a 6-point factuality scale, ranging from Very Low to Very High. For each post in our datasets, we systematically extract, expand, and parse all embedded URLs, checking whether the URL belongs to a low- or high-credibility domain.
@@
Given recent news777www.bbc.com/news/articles/c8rx28v1vpro888www.state.gov/u-s-department-of-state-takes-actions-to-counter-russian-influence-and-interference-in-u-s-elections/ and prior instances of interference by Russian agencies in U.S. elections (Badawy et al., 2019), we also examined the potential amplification of Russian media outlets. Following a similar approach to previous work (Pierri et al., 2023), we utilize the VoynaSlov dataset (Park et al., 2022) to obtain a list of 23 state-affiliated Russian websites. We then check whether the extracted URLs link to one of these web domains.
@@
Finally, we employ a similar approach to identify content linked to conspiracies, such as QAnon, given the prevalence of fringe and conspiratorial theories during the 2020 US Election, and its aftermath (Sharma et al., 2022; Vishnuprasad et al., 2024). Building on the methodology of (Wang et al., 2023), we detect posts sharing QAnon content by utilizing a list of keywords commonly associated with the conspiracy, as outlined by (Sharma et al., 2022).
5. Results
@@
This section presents the results obtained from the detection of online coordination [^9] Code publicly available at github.com/mminici/USElection-CoIA-TheWebConf25. We first present the results for the detection of web domain promotion and content amplification. Following the identification of these campaigns, we examine the content pushed by coordinated networks and the engagement received by suspicious actors focusing on several interaction and engagement metrics.
5.1. Detecting CoIA for Web Domain Promotion (RQ1)
@@
We detect CoIAs that push web domains within and across platforms as described in Section 4.1.1. We begin by presenting results for each coordinated network within individual platforms. Following this, we analyze the inter-platform CoIA.
@@
Figure 1. Cross-platform coordination network illustrating user coordination across the three social media platforms. The percentages at the bottom right indicate the proportion of users from each platform within this network.
5.1.1. Intra-Platform Coordination.
We first analyze each intra-platform CoIA based on the co-URL similarity networks extracted from each platform separately.
Telegram: We identified 33 highly coordinated channels co-sharing URLs to web domains with a partisan slant and low factuality. As shown in the Table 2, the most frequently shared domains are predominantly right-leaning and of low credibility, as assessed by MBFC101010mediabiasfactcheck.com/ and consistent with previous findings (Biswas et al., 2024). Notably, the table highlights the presence of a Russian state-controlled media outlet (RT.com) and its video-on-demand subsidiary (Ruptly.tv), which could indicate a foreign effort to interfere in the election. Both RT.com and Ruptly.tv have previously been accused of orchestrating campaigns to influence U.S. Elections via social media111111www.nbcnews.com/politics/2020-election/facebook-blocks-russia-backed-accounts-other-sites-keep-churning-out-n1242683.
Additionally, this CoIA promotes a far-right website (thegatewaypundit.com) and an extremist-friendly video platform (odysee.com), suggesting ties to fringe ideas. A manual review of these Telegram channels, including their profile descriptions (see Table D.3 in the Appendix) and shared messages (see Table D.7 in the Appendix), reveals a strong prevalence of content and accounts promoting conspiracy theories, particularly those related to COVID vaccines, 5G, and alternative news media.
@@
Table 2. Top-5 domains shared in Telegram by coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on MBFC.
$\mathbb{X}$/Twitter: We identified a network of 19 coordinated accounts predominantly promoting content from right-leaning domains, as listed in Table 3. A manual review of this $\mathbb{X}$ CoIA reveals that many of these coordinated accounts share similar profile descriptions, reflecting narratives tied to religious and conservative principles (see Table D.4 in the Appendix). Additionally, an analysis of the messages (see Table D.8 in the Appendix) shows a strong presence of politically partisan content, predominantly aligned with right-leaning ideologies.
@@
tmz.com 5 Mixed RIGHT-CENTER
conservativeinstitute.com 5 Mostly RIGHT
@@
Table 3. Top-5 domains shared in $\mathbb{X}$ by coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on MBFC.
@@
Facebook: On Facebook, no well-known media outlets dominate the shared content. However, the most shared domain is a partisan news outlet, as suggested by the semantics of its title: gorightnews.com. Additionally, there is a direct connection to a public activist known for political campaigns121212en.wikipedia.org/wiki/Peter_Boykin and their website: peterboykin.com. These campaigns and related Facebook accounts primarily focus on the “Gays for Trump” movement, an American LGBTQ organization that advocates for U.S. President Donald Trump131313en.wikipedia.org/wiki/Gays_for_Trump. The top-engagement messages (see Table D.9 in the Appendix) and bios (see Table D.5 in the Appendix) indicate strong support for right-leaning ideologies.
5.1.2. Cross-Platform Coordination.
@@
The analysis of the cross-platform network reveals distinct patterns. The unified cross-platform network, displayed in Fig. 1, highlights a giant component connecting coordinated accounts from both Telegram and $\mathbb{X}$, while the Facebook coordination network remains largely disconnected. A few noteworthy observations:
@@
First, we observe that most cross-platform connections occur between $\mathbb{X}$ accounts and Telegram channels. Examining the biography descriptions of the top users by degree (see Table D.10 in the Appendix), we note a prominent presence of non-mainstream news outlets and conspiracy theories, such as the flat earth one.
@@
Second, examining bridge nodes between the Telegram-coordinated cluster and the cross-platform giant component, we find highly influential channels with up to 24,000 subscribers. These channels often feature bios referencing free speech and religion, such as:
  • The FIGHT for FREEDOM - :Aron TRUTH Social
  • Q Reee-searchers Watchers
@@
  • ‘Arise and shine, for your light has come, and the glory of the LORD rises upon you.’ Isa. 60:1 WE RISE beyond the challenges of yesterday to become better stewards of tomorrow!
@@
Similarly, the bridge node between the $\mathbb{X}$-coordinated cluster and the cross-platform giant component is a highly influential account with 32,000 followers. The bio reads: “NO DM’S. Beautiful disaster. Self-proclaimed arbiter of great ideas. Here to annoy the dumb asses. NO LISTS!! #Imvotingforafelon #animallover”, indicating a strong partisan flavor.
@@
Finally, the largest component of the cross-platform coordination network is dominated by a mix of Telegram and $\mathbb{X}$ accounts promoting domains such as magapac.com, QAnon-related narratives (e.g., “WWG1WGA”141414en.wiktionary.org/wiki/WWG1WGA), and accounts with partisan bios like:
@@
  • We The Ultra Patriots is made up of patriotic Americans dedicated to exposing crimes against humanity, false flags, lies, and corruption. We The Ultra Patriots are…
@@
A closer look at the top shared domains within the coordinated network (see Table 4) reveals that most are partisan, including truthsocial.com, an alternative non-mainstream social media platform151515en.wikipedia.org/wiki/Truth_Social. Additionally, six of the shared domains are Russian state-affiliated websites161616github.com/chan0park/VoynaSlov/tree/master, including ruptly.tv (2,117 shares), rbc (8 shares), gazeta (7 shares), ruptly.video (2 shares), tv5 (1 share), and mil (1 share), further suggesting potential Russian information operations to interfere with the election discourse. Notably, coordinated actors, just 1% of observed Telegram accounts, shared over 2,000 messages linking to these media outlets—nearly 50% of all links—demonstrating their outsized role in content amplification and none of these accounts is linked to any media organization. Examining the posts and messages with the highest engagement (see Table D.10 in the Appendix), we observe the prevalence of ultra-MAGA narratives, conservative religious themes, and environmental news.
@@
Table 4. Top domains shared by cross-platform coordinated accounts. Columns represent, from left to right: domain, the number of shares among coordinated accounts, and the factuality and political leaning scores based on MBFC.
@@
Figure 2. Content analysis of coordinated accounts on Telegram, $\mathbb{X}$, and Facebook. We report the top-10 most recurring topics and compare their prevalence against the organic discourse.
5.2. Detecting CoIA for Content Amplification (RQ2)
In this section, we present our findings on the specific topics amplified by coordinated users within content amplification campaigns. Following the method in §4.1.2, we extract the Text Similarity Network and identify a seed set of coordinated users. We then apply BERTopic, as described in §4.2.1, to identify themes these users aim to promote and compare them to themes that emerge organically on each platform. For each platform, we report the top-10 most prevalent topics within the coordinated cohort and compare their prevalence to that within the organic population. This approach highlights themes promoted by coordinated users, with large differences indicating topics characteristic of coordinated activity. In the remainder of this section, we detail our findings on coordination-driven content amplification across the three platforms.
Telegram: Based on the significantly high similarity of their shared content, we identified 57 coordinated Telegram channels. Among these channels, public health is a prominent topic of discussion, as can be seen in Figure 2(a), with skepticism around the COVID-19 vaccine emerging as the most frequently discussed theme. Additionally, many messages promote alternative medicine, and there is notable discussion surrounding a conspiracy theory related to bird flu vaccination. Telegram’s coordinated channels also focus on strictly political discussions, particularly on the U.S. Presidential debate, the Russia-Ukraine conflict, and Donald Trump’s legal issues. Interestingly, we observe discourse around immigration issues in Ireland, which appears to be echoed by accounts associated with the MAGA movement. Notably, three out of the ten topics involve conspiracy theories related to flat-earth beliefs, bird flu, and antisemitism (see Table E.12 in the Appendix).
We also assess the presence of QAnon-related keywords in content shared by both coordinated and organic channels. Approximately 1.66% of content from coordinated users contains at least one QAnon-related keyword, compared to only 0.12% of content from the organic population. This substantial difference is evident in Table 5, where we list the top five keywords present in the coordinated group, also compared to keywords shared by organic channels. Although the 57 coordinated channels make up only 0.36% of the total number of Telegram channels in our dataset, they show a marked prominence, both in absolute and relative terms, in sharing QAnon-related keywords compared to organic channels. These findings indicate a propensity for Telegram CoIAs to disseminate conspiracies.